### v2

# Bridging the Culture Gap: A Framework for LLM-Driven Socio-Cultural Localization of Math Word Problems in Low-Resource Languages

Abstract
Large language models (LLMs) have demonstrated significant capabilities in solving mathematical problems expressed in natural language. However, multilingual and culturally-grounded mathematical reasoning in low-resource languages lags behind English due to the scarcity of socio-cultural task datasets that reflect accurate native entities such as person names, organization names, and currencies. Existing multilingual benchmarks are predominantly produced via translation and typically retain English-centric entities, owing to the high cost associated with human annotater-based localization. Moreover, automated localization tools are limited, and hence, truly localized datasets remain scarce. To bridge this gap, we introduce a framework for LLM-driven cultural localization of math word problems that automatically constructs datasets with native names, organizations, and currencies from existing sources. We find that translated benchmarks can obscure true multilingual math ability under appropriate socio-cultural contexts. Through extensive experiments, we also show that our framework can help mitigate English-centric entity bias and improves robustness when native entities are introduced across various languages.
Bridging the Culture Gap: A Framework for LLM-Driven Socio-Cultural Localization of Math Word Problems in Low-Resource Languages
Israel Abebe Azime1,∗, Tadesse Destaw Belay2,∗, Dietrich Klakow1 Philipp Slusallek1, Anshuman Chhabra3 1 Saarland University, Saarland Informatics Campus 2 Instituto Politécnico Nacional, 3 University of South Florida, Tampa, FL USA
Equal Contribution.
1 Introduction
Figure 1: An example showcasing an English math word problem, its direct translation, and an automatically localized version with culturally adapted entities. While the problem structure remains identical, large language models (LLMs) often fail to answer correctly when entity names or currencies are altered. This highlights a key limitation in current LLM robustness. In this paper, our goal is to audit models and rectify their robustness issues so that remain consistent and accurate across such culturally grounded variations.
Mathematical reasoning has been adopted as a core milestone in large language model understanding research Yan et al. (2024); Ahn et al. (2024a, b). Math word problems (MWPs), a key component of mathematical reasoning tasks, have been widely explored as a challenging benchmark for LLMs Srivatsa and Kochmar (2024). MWPs are characterized by their integration of mathematical knowledge with scenarios drawn from everyday activities, which can vary across cultures. The ability of low-resource and multilingual LLMs to solve MWPs correctly depends on multiple factors: (1) the language used to prompt the models Adelani et al. (2024) and (2) the linguistic complexity of the questions Srivatsa and Kochmar (2024). Despite their impressive performance on MWPs, a crucial question remains: do LLMs truly retain performance under culturally diverse MWPs, even when the underlying mathematical structure remains unchanged?
Low-resource language (LRL) math word problem evaluation heavily relies on human or machine-translated benchmarks, as demonstrated by the translation of GSM8K Cobbe et al. (2021) (a dataset of 8.79K high-quality, linguistically diverse grade school math word problems) into GSM8K_zh for Chinese Yu et al. (2023), MGSM for ten typologically diverse languages Shi et al. (2022a), and AfriMGSM for 18 African languages Adelani et al. (2024). Existing human- or machine-translated benchmarks often fail to account for local cultural contexts that contain day to day activities, common names and currencies. As a result, evaluating LLMs on such non-localized data primarily measures their mathematical understanding ability in English centric scenarios, names and currencies.
In this work, we investigate the task of creating more culturally aligned math word problems from their translated variants. Owing to the high costs associated with human-based localization, it is imperative to develop automated frameworks capable of producing large-scale datasets. Therefore, this motivates our first research question in this work: (RQ1) Can LLM-driven pipelines be used to localize translated benchmarks for low-resource languages?
Moreover, given that a significant number of benchmark datasets are created using translation  Alabi et al. (2025) and do not study appropriate socio-culturally grounded datasets (see Figure 1) we propose an automated framework for socio-cultural localization of MWPs in low-resource languages. Using our framework, we can study our second research question, as follows: (RQ2) Does introducing socio-cultural local entities in translated existing benchmarks reveal performance disparities? We undertake several experiments to study RQ2 for both low-resource languages (LRLs) and English to ascertain the performance gaps. Finally, we utilize our framework to generate localized MWPs for 18 African Languages covered by Adelani et al. (2024) and investigate our final research question: (RQ3) Can automated localization improve model robustness by augmenting benchmarks with culturally adapted variants?
To address these questions, our work makes the following contributions:
  • We develop an LLM-driven localization pipeline to generate culturally adapted versions of translated datasets by replacing key entities with apt socio-cultural variants. [^1] https://huggingface.co/collections/israel/cultural-localilzation-687e392f6be2aaa0b7c7831c
  • Further, through extensive experiments enabled via our framework, we investigate performance disparities between localized and directly translated benchmarks across several LLMs like Gemma and LLaMA models. Our findings reveal how the presence of cultural entities influences LLMs’ ability to solve mathematical word problems. For instance, we observe as much as 9% (numeric match) drop in performance for GPT-4o-mini  (with similar trends seen across other LLMs).
  • Then, using the localized data generated by our automated localization framework, we fine-tune LLMs on these socio-cultural MWP variants and find that model robustness and generalization improves significantly across multiple languages, thereby paving the way for improving the performance of LLMs in a straightforward manner.
Stages Example
1 English $x_{\text{en}}$ Mandy owes Benedict \$ 100. They agreed to have monthly interest of 2%. If Mandy was able to pay it after 3 months, how much should she give to Benedict?
2 Translated
$x_{\text{trans}}\leftarrow x_{\text{en}}$ Mandy anadaiwa \$ 100 na Benedict . Wamekubali kuwa na riba ya kila mwezi ya 2%. Ikiwa Mandy aliweza kulipa baada ya miezi 3, anafaa kumpa Benedict pesa ngapi?
3 Important entities
[…]
4 replacement dictionary
replacement_dict=fn(en,db)replacement\_dict=fn(en,db) {‘‘Mandy’’: ‘‘Camari’’, ‘‘Benedict’’: ‘‘Julani’’, ‘‘\$’’: ‘‘shilingi’’, ‘‘dollar’’: ‘‘shilingi’’}
5 Entity replaced
$x_{\text{ent}}=fn(x_{\text{ent}},replacement\_dict)$ Camari owes Julani shilingi 100. They agreed to have monthly interest of 2%. If Camari was able to pay it after 3 months, how much should she give to Julani?
6 Auto Localized
[…]
7 Quality Check
length($\hat{x}_{\text{loc}}$) == length($x_{\text{trans}}$), key entities not in $\hat{x}_{\text{loc}}$, replacement entities in $\hat{x}_{\text{loc}}$ , and similarity($\hat{x}_{\text{loc}}$, $x_{\text{trans}}$) > 0.8, Full human verification for the test set and sampled human verification for the training set. if failed return ($x_{\text{trans}}$)
Table 1: The different stages of our MWP automated localization framework for low-resource languages. We show a step-by-step transformation from English to a culturally and linguistically localized Swahili version: direct translation, name and currency replacements, a semi-localized substitution, and the final fluent localization. Colored highlights indicate aligned entities across languages; occurrences of LLM denote the use of an LLM.
2 Related Work
Math World Problems.
Math word problems are mathematical questions framed in everyday language, often grounded in real-life scenarios and activities rather than expressed purely through abstract symbols or equations Cobbe et al. (2021). These problems serve as an engaging way to learn mathematical reasoning, as they require not only computational skills but also the ability to interpret and model real-world situations.
LLM Robustness.
Although LLMs are highly capable of solving complex tasks, they often struggle with simple variations in input (Askari et al., 2025; Chhabra et al., 2024; Achara and Chhabra, 2025). For example, Abedin et al. (2025) demonstrate that even minor perturbations such as spelling mistakes can significantly impact model performance, a challenge that is particularly pronounced in low-resource settings and evaluations.
Translation and Localization.
Translating existing benchmarks is one of the simplest approaches to benchmark creation, particularly for low-resource languages that lack dedicated evaluation datasets. In such cases, researchers often resort to translating established benchmarks from high-resource languages to enable evaluation and comparison Adelani et al. (2024); Koto et al. (2024); Li et al. (2023); Son et al. (2024). In African context close to 30% of resource papers translate existing benchmarks to create language specific benchmarks Alabi et al. (2025). However, to better reflect native cultural identities and move away from the Western-centric concepts that often persist in translated benchmarks, some researchers have begun developing fully localized benchmarks Yu et al. (2025). This work complements manual localization and cultural adaptation efforts by reducing the scale and resources required to acquire such data. Just as traditional data augmentation methods improve model robustness without the need for entirely new data creation, our proposed approach similarly complements translation by enabling the creation of large-scale, culturally adapted datasets with minimal overhead.
Automatic Entity based Augmentation.
In addition to manual localization efforts carried out by volunteers, Ye et al. (2024) introduce a novel data augmentation technique that leverages large language models (LLMs). Specifically, they apply this method to enhance performance in few-shot Named Entity Recognition (NER) tasks, demonstrating that LLM-driven augmentations can serve as a valuable complement to human-curated resources.
Mathematics and Cultural Entities.
Karim et al. (2025); Tomar et al. (2025) examined the influence of cultural context on mathematical problems by analyzing the impact of culturally grounded entities such as personal names and food items. However, one critical dimension that remains underexplored is the role of language itself and the biases that can be propagated through translation. Since the significant number of multilingual training and evaluation datasets are created via translation from high-resource languages Alabi et al. (2025), they often fail to capture the linguistic and cultural nuances of the target languages. Evaluating model robustness in the presence of culturally specific entities is a critical challenge, especially as LLMs are increasingly deployed in real-world applications. Native language speakers naturally refer to familiar names, organizations, and currencies from their own cultural context, making it essential for models to handle such variations reliably.
In this work, we address this gap by developing an automated pipeline for creating localized datasets in diverse languages, studying how language and culturally specific entities affect model performance, improving robustness to entity-level variations through fine-tuning, and evaluating which data creation strategies best support both scalable generalization and cultural alignment.
3 Methodology
3.1 Pipeline
Figure 1, shows our proposed automated pipeline for socio-cultural localization on MWPs. We first extract personal names, organization names, and currencies, and then replace them with manually collected local entities. Then, we generate accurately localized training and test sets for our experiments. Unlike LLM-based localization, our approach ensures that all relevant entities, particularly those critical to the problem, are consistently and correctly replaced. Additionally, we incorporate manual verification steps to further improve localization quality and ensure high fidelity in culturally adapting the benchmarks.
Below, we outline the key stages of our framework pipeline and discuss the design choices that guided its development:
Entity Classification. Our pipeline processes each word in the input text and classifies it as a personal name, organization name, or currency. We chose to focus on these entity types to enable controlled generation and replacement, ensuring that the original meaning of the sentence is preserved and that no unintended or confusing content is introduced. Replacing animal and food names tend to generate sentences that lack contextual meaning even though our pipeline can handle it easily. While we evaluated several Named Entity Recognition (NER) methods Tjong Kim Sang and De Meulder (2003) and POS tagging using spaCy [^2] https://spacy.io/usage/linguistic-features, we found that large language models (LLMs) with structured output formats were significantly more robust. In particular, they handled variations in spelling and casing more effectively, which are common failure points for traditional models. Additionally, using LLMs for entity classification enables a more scalable pipeline, allowing for easy extension to include additional entity types based on the requirements of the task. Furthermore, we conducted most of the processing in English, as we observed that English entities are more reliably identified when operating in the English language.
Multilingual Entity Database. To ensure that the entities used for replacement were culturally relevant, we collaborated with a team of volunteers to curate unisex personal names, organization names, and representative currency values for each language. Special attention was given to selecting unisex names to avoid introducing gender-specific biases or incorrect pronoun associations.
Replacement Dictionary Creation. Using the output of the entity classification step, we assign replacement entities to each extracted item by referencing a multilingual entity database, as illustrated in Stage 4 of Table 1.
Auto Localization. Our pipeline operates on three versions of each input sample: the original English sentence ($x_{\text{en}}$), its direct translation into a target language ($x_{\text{trans}}$), and an entity-replaced version of the translation ($x_{\text{ent}}$), as illustrated in Table 1.
To generate a properly localized version of $x_{\text{ent}}$, we use a one-shot prompting setup. Specifically, we construct a prompt by showing the LLM the pair $(x_{\text{en}},x_{\text{trans}})$ as an example, and then ask it to translate $x_{\text{ent}}$ accordingly:
$\text{LLM}\left(\left[(x_{\text{en}},x_{\text{trans}}),\ x_{\text{ent}}\right]\right)\rightarrow\hat{x}_{\text{loc}}$
This method provides stronger contextual grounding and helps the model preserve the structure and fluency of the target language. Empirically, we found it to be more effective than directly prompting the LLM to localize text without reference examples.
Quality Check Blocks. Between most of the modules in our pipeline, we applied lightweight quality checks to ensure accurate and consistent localization. The core motivation for building this controlled pipeline, as opposed to relying solely on fully automated localization with LLMs, is to guarantee that the model either produces a localized version of the text or returns the original non-localized text if no entities are detected. This conditional behavior prevents unnecessary modifications and maintains data integrity.
As shown in stage 7 of Table 1, our quality control measures ensure no overlapping or inconsistent entity replacements, enforce a single currency type per problem to avoid conversion errors, verify that localized outputs match the length of their direct translations, and check similarity between localized and translated text using the difflib library333docs.python.org/3/library/difflib.html. Since all languages share the same entity-replaced text ($x_{\text{ent}}$), human verification only requires comparing entities in the replacement dictionary. This prevents unnecessary additions by LLMs, maintains prompt consistency, and keeps outputs close to the original structure.
LLMs for Localization. In this work, we leveraged Gemini models Comanici et al. (2025) due to their exceptional multilingual performance shown by natural language generation (NLG) multilingual benchmarks Ojo et al. (2023). To reduce experimental costs while assessing our approach, we used Gemini-1.5-pro for evaluation of the framework, and employed Gemini-2.5-pro for the final data generation.
3.2 Datasets Used
Evaluation Dataset. AfriMGSM Adelani et al. (2024) is a manually translated benchmark spanning 18 languages sourced from MGSM Shi et al. (2022b). We found that approximately 86% of the test set, includes at least one important entity. Using the English and manually translated pairs, we created a localized evaluation dataset by replacing these entities with culturally appropriate local alternatives while keeping the rest of the problem content unchanged. The correctness of the localized dataset was verified through manual inspection by the authors. Since the approach returns the unlocalized item if it fails, no additional noise is introduced.
Figure 2: Human Validated Comparison of Localization Quality Between Auto Localization and Direct Prompting (Gemini-1.5-pro). Our auto localization framework produces significantly better and appropriate culturally localized outputs compared to direct prompting, which often fails to adapt entities to the target culture. This highlights the advantage of our method in achieving consistent and controllable localization.
Table 2: Datasets used in our auto localization framework and their details (sources, size, split).
source split #source #localized #lang.
GSM8K train 8790 25100 18
AfriMGSM test 4500 4500 18
Training Dataset. Due to the lack of manually translated training datasets for math word problems across all target languages, we leveraged the GSM8K dataset from Cobbe et al. (2021) and applied our pipeline to generate translated and localized versions. For translation, we used the NLLB-200-3.3B model NLLB Team et al. (2022), which we consider to offer the best trade-off between model size and translation quality among open-source models that support all the languages considered in this work.
To ensure translation quality, we filtered the outputs using SSA-COMMET Li et al. (2025), a sentence-level semantic similarity metric that scores translation quality on a scale from 0 to 1, with 1 indicating perfect translation. Based on the score distribution shown in the Appendix D, we retained only translations with a COMMET score above 0.65. From these, we selected the top 1,500 examples per language for automatic localization and use as training data in our after localizing each. While intermediate experiments leverage Gemini-1.5-pro to evaluate the pipeline and assess its effectiveness due to cost constraints and to avoid unwanted behaviors, such as prompt caching [^4] https://ai.google.dev/gemini-api/docs/caching?lang=python, the final dataset is generated using Gemini-2.5-pro to ensure the highest quality standards.
Figure 3: Direct Translation vs. Auto Localization performance for Aya-expanse-32b (top left), LLaMA-3-70B-Instruct (top right), Gemma-2-27b-it (bottom left), and GPT-4o-mini (bottom right). We observe that some languages, such as Swahili, are more prone to confusion from newly introduced entities in three out of four models, highlighting the lack of robustness for local entity presence in low-resource languages.
Figure 4: Impact of Cultural Entities on English benchmarks. Does the presence of cultural entities ($x_{\text{ent}}$; see Table 1) affect model performance? Results are shown for Aya-expanse-32b (top-left), LLaMA-3-70B-Instruct (top-right), Gemma-2-27b-it (bottom-left), and GPT-4o-mini (bottom-right). We observe that the presence of local entities has a consistent effect on evaluations, with the exception of Gemma-2-27b-it.
4 Experimental Setup
Evaluated LLMs. In this work, we evaluated both open-source and closed-source models commonly studied in existing multilingual mathematical research. Our selection spans a range of model sizes, including Aya-expanse-32b Dang et al. (2024), LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct Grattafiori et al. (2024), Gemma-2-9b-it, Gemma-2-27b-it Team et al. (2024) and GPT-4o-mini. The finetuning experiments are done only on LLaMA-3-8B-Instruct, Gemma-2-9b-it because of compute constraints.
Evaluation Metrics. Most existing work on mathematical word problems focuses on evaluating reasoning capabilities, where intermediate calculation steps are used as part of the input or as supervision for general benchmarking. However, due to the lack of multilingual reasoning benchmarks and the challenges involved in extracting step-by-step reasoning across 18 different languages, this research focuses solely on evaluating final answers.
Several metrics are commonly used in mathematical evaluation research, including exact match, numeric match. In this work, we select the metrics that best align with the core objectives of our study. Exact Match (EM) evaluates whether the predicted answer exactly matches the reference answer as a string. Numeric Match (NM) checks whether the predicted numerical value matches the ground truth, ignoring differences in formatting such as units or punctuation after converting them into floating point data and account for errors between them. Moreover, for ensuring robustness and minimal noise via prompt selection, we adopted three prompt variations from Adelani et al. (2024) for our experiments. We further customized them to return only the final output without intermediate steps and reported the average score across all three prompts.
Reproducibility. We used ElutherAI’s open source Language Model Evaluation Harness (lm-eval) framework Gao et al. (2024) to evaluate models. Instead of direct model querying this allows us to have standardized reproducible evaluations across all LLMs. We also set generation parameters (i.e. temperature) to zero for consistency.
Language Translated English with Entities Auto Localized All Data
$x_{trans}$ $x_{ent}$ $\hat{x}_{\text{loc}}$
[…]
# Native Robust lang. 6 4 7 11
Language Translated English with Entities Auto Localized All Data
$x_{trans}$ $x_{ent}$ $\hat{x}_{\text{loc}}$
[…]
# Native Robust lang. 7 8 12 10
Figure 5: Native Robustness for Model (Numeric Match $\Delta$). For each language, we report $\Delta_{\mathrm{NM}}=\mathrm{NM}_{\text{localized}}-\mathrm{NM}_{\text{translated}}$ after LoRA fine-tuning of LLaMA-3-8B-Instruct (left) and Gemma-2-9b-it (right) across training-data configurations. Positive denotes higher accuracy on localized benchmarks (greater robustness to local entities); Negative denotes higher accuracy on English-centric benchmarks (reduced robustness to localization) and Yellow denotes that the model performance did not degrade or improve. Languages are ordered from lower- to higher-resource by FineWeb2 coverage Penedo et al. (2025).
5 Results and Analysis
We now present the results of our experiments across several LLMs and our generated localized datasets to answer each of our RQs:
(RQ1) Can LLM-driven pipelines be used to localize translated benchmarks?
To address the high cost associated with using human translators for creating localized versions of datasets, we introduce an automatic pipeline that generates culturally adapted versions of mathematical word problems. While this pipeline is not intended to replace human annotators, it serves as a valuable tool for producing augmented datasets that help improve the robustness of LLMs to entity-based variations in problems.
In our first set of experiments (please see Figure 2), we used Gemini-1.5-pro to generate localized versions of the dataset using both our six-stage localization pipeline and direct prompting via the Gemini API. Once the two versions of the localized dataset were generated, three annotators independently labeled each instance as either a valid or invalid localization. Additionally we wanted to show this method does not require the latest/largest models – allowing us to achieve stellar performance even with older models like Gemini-1.5-pro.
Next, we compare final downstream task performance across directly translated and auto localized versions of MWPs. These results are provided in Figure 2 and illustrate the effectiveness of our localization pipeline compared to manual prompting. In addition to improved performance and more accurate localization, our method offers greater control over the types of modifications applied. Manual prompting performs well when there are direct equivalents for names across languages, but it often returns the original translation rather than a properly localized version. This limits its effectiveness in producing culturally adapted outputs. In Figure 2 items that are not localized include, names or organization names the pipeline was not able to capture and fallback output because quality checker.
(RQ2) Does introducing local entities in translated benchmarks reveal performance disparities for MWPs in LRLs and English?
We first analyze LRLs and then English language MWPs. Given that 30% of newly created African datasets are based on translated content Alabi et al. (2025), it is important to assess whether LLMs are overfitting to English-centric entities commonly preserved during translation. Such overfitting may lead to performance degradation when these entities are replaced with their native counterparts. In this work, we hypothesize that models should perform better when evaluated on data containing entities from their target (native) language, as opposed to data where entities originate from other languages.
Figure 3 illustrates the performance differences between models evaluated on translated benchmarks and those evaluated on automatically localized benchmarks, highlighting the impact of entity localization on model robustness. We can observe that the Aya-expanse-32b LLM performs well on translated benchmarks while LLaMA-3-70B-Instruct attains a performance gain from seeing native entities. Gemma-2-27b-it has more consistent results across native and English centric entities.
Next, Figure 4 shows the performance of math word problems in English where personal names, organization names, and currencies have been replaced with entities from the respective target languages. We compare these results with baseline scores obtained from the original English benchmark, which contains English-centric names and currencies. This comparison allows us to evaluate the impact of introducing culturally specific entities into English problem statements.
Both LLaMA-3-70B-Instruct and GPT-4o-mini achieve higher accuracy on the localized (native-entity) variant than on the English benchmark, whereas the Aya-expanse-32b model’s performance lags behind. Gemma-2-27b-it generally tracks the localized variant more closely, despite occasional dips and overshoots.
(RQ3) Can automatic localization improve model robustness by augmenting benchmarks with culturally adapted variants?
Figure 5 presents the performance of models trained on different flavors of MWP data. We begin with models trained on the translated dataset, followed by those trained on the original English GSM8K dataset, and then models trained using the training data produced by our localization pipeline. Finally, we include results for models trained on a combined dataset that includes all three sources.
Our evaluation compares model performance on the localized benchmark versus the translated benchmark. This comparison helps us understand whether the inclusion of native cultural entities affects model behavior. A positive difference indicates that the model performs better on the localized benchmark, suggesting improved robustness to cultural variations. Conversely, a negative difference highlights cases where the model handles English-specific entities such as names and currencies effectively but fails when presented with culture-specific alternatives.
We observe that both LLaMA-3-8B-Instruct and Gemma-2-9b-it models exhibit improved robustness to entity changes in several languages. For LLaMA-3-8B-Instruct, the best performance is achieved when localization is combined with additional noisy data, indicating that diverse training sources can enhance generalization. In contrast, for Gemma-2-9b-it, models trained exclusively on Auto Localization data perform best, suggesting that this model benefits more from clean, culturally adapted input.
Additionally, we find that languages with more training resources tend to benefit more from localization. This trend is evident from the table, where languages are sorted in ascending order based on their availability in the FineWeb2 Penedo et al. (2025) dataset. Higher-resource languages (appearing lower in the table) show greater performance gains, while lower-resource languages (toward the top) experience more limited improvements. More exploration to see characteristics of languages affected by the finetuning experiment included in the Appendix C.
6 Conclusion
Due to the scarcity of native, low-resource mathematical reasoning datasets that include local entities, translation remains the predominant source of benchmark questions. However, performance on these translated benchmarks is highly sensitive to English-centric terms. We present a framework that culturally localizes translated datasets into variants enriched with local entities . We highlight the biases and instabilities introduced by translation-only benchmarks and show that our localization framework improves model robustness in the presence of native entities.
Greater emphasis should be placed on creating MWPs that center around the cultural activities of the target community, in addition to incorporating cultural names into existing ones. For future work, we will use this framework for more reasoning-focused evaluations, create strong multilingual training datasets, and extend the approach beyond mathematics to complement translation-based benchmarking in other tasks.
Limitations
Due to the high cost associated with human-centered localization, we developed an automatic localization pipeline capable of generating culturally relevant datasets from translated, English-centric data. While this pipeline offers a scalable and efficient solution for large-scale dataset creation, it is not intended to replace human annotation. In scenarios where the cost of data acquisition is not a constraint, human-centered localization should be preferred for its higher accuracy and cultural fidelity. The primary advantage of our automated approach lies in its ability to support the expansion of training pipelines across multiple languages and domains with minimal manual effort. We galvanize the community to place greater emphasis on developing math word problems rooted in everyday community activities, creating appropriate socio-cultural scenarios that can be used to further improve models.
Acknowledgment
The authors would like to thank the German Federal Ministry of Education and Research and the German federal states (http://www.nhr-verein.de/en/our-partners) for supporting this work/project as part of the National High-Performance Computing (NHR) joint funding program. We also thank PaliGemma academic program for gemini credits used in this research.
Appendix
Appendix A Prompt Templates Used
Below, we present the prompts used for our Auto-Localizer and direct prompt localization. We observed that including the intermediate steps shown in Table 1 improves the accuracy of localization.
Prompt: Entity Localization Task You are an expert linguistic assistant. Your task is to edit a sentence in a native language to match a change made in its English parallel.
Here is the context: Original English: The original sentence. Original Native: The original translation of the English sentence in {native_lang}. Modified English: The English sentence has been edited. One or more words have been replaced. Your goal is to produce a Modified Native sentence by applying the exact same replacement to the Original Native sentence.
Crucial Instructions: DO NOT re-translate the entire sentence. Only replace the specific words that were changed in the English version. Preserve the original grammar and structure of the native sentence as much as possible. Ensure the final Modified Native sentence is natural and grammatically correct in {native_lang}. Respond with ONLY the Modified Native sentence and nothing else.
Example: Original English: Janet’s ducks lay 16 eggs per day. Original Native (French): Les canards de Janet pondent 16 œufs par jour. Modified English: Andrea’s ducks lay 16 eggs per day. Modified Native (French): Les canards d’Andrea pondent 16 œufs par jour.
Your Task: Original English: {original_eng} Original Native ({native_lang}): {original_native} Modified English: {modified_eng} Modified Native ({native_lang}):
Prompt: Direct Localization using LLMs without our pipeline You are an expert linguistic assistant specializing in localization. Your task is to localize the following English sentence into {target_lang_name}.
Localization means adapting the text to the target culture and context by: Identifying and replacing English-specific entities (such as person names, organization names, and currency symbols/names) with culturally appropriate equivalents in {target_lang_name}. Ensuring the localized text sounds natural and grammatically correct in {target_lang_name}. Crucially, the localized text should be generated directly from the English sentence, without relying on any pre-existing native translation. English Sentence to Localize: {original_eng} Your Goal: Produce a Localized {target_lang_name} sentence. Instructions: Identify person names, organization names, and currency symbols/names in the English sentence. Replace these entities with culturally appropriate {target_lang_name} alternatives. If an entity does not have a direct cultural equivalent or is already culturally neutral, it may remain unchanged. Respond with ONLY the Localized {target_lang_name} sentence and nothing else. Localized {target_lang_name}:
Appendix B Additional Evaluation Results
Figure 6 presents additional evaluations, extending the analysis to a broader set of models with diverse characteristics.
Figure 6: Translated vs Auto Localization performance Gemma-2-2b-it (top left) Llama-3.2-1B-Instruct (top right) Qwen3-4B (bottom left) LLaMA-3-8B-Instruct (bottom right)
Appendix C Precence of Cultural entities in our Training data
Figure 7 illustrates the distribution of culturally relevant items within the 1,500 data samples selected for each language. Our analysis reveals that a considerable proportion of the translated items could not be localized to the target language, primarily because they did not contain the types of entities, such as personal names, organizations, or currencies, that form the basis of our localization strategy. This highlights a key limitation of translation-only benchmarks: when source items lack culturally grounded entities, they fail to provide meaningful opportunities for localization, thereby reducing the effectiveness of culturally adapted datasets. By quantifying this gap, we emphasize the importance of designing datasets that explicitly incorporate entity-level cultural relevance to ensure robust evaluation and training across languages.
Figure 7: Number of Data Samples Without Cultural Entity Replacements out of 1500 selected, Grouped by Language for Training dataset.
Appendix D Translation quality - SSA-COMMET score
Figure 8: Comet score caption here, Score ranges from 0-1, 1 is high translation quality
Appendix E Code and Reproducibility
We used ElutherAI’s open source Language Model Evaluation Harness (lm-eval) framework Gao et al. (2024) to evaluate models. Instead of direct model querying this allows us to have standardized reproducible evaluations across all LLMs. We also set generation parameters (i.e. temperature) to zero for consistency.